An electric potential field is produced by point charges 1 mC and 4 mC located at (-2,1,5) and (1,3,-1) respectively. The energy stored in the field is


1
Expert's answer
2021-07-04T17:40:58-0400

"q_1=1mc;q_2=4mc"

Point A(-2,1,5); point B(2,3,-1)


"r_{AB}=\sqrt{{(x_2-x_1)^2}+(y_2-y_1)^2+(z_2-z_1)^2}"

Put value


"r_{AB}=\sqrt{{(1+2)^2}+(3-1)^2+(-1-5)^2}"

"r_{AB}=7"


"U=\frac{kq_1q_2}{r_{AB}}=\frac{9\times10^9\times1\times10^{-3}\times4\times10^{-3}}{7^2}=734J"